OpenStudy (anonymous):

find the probability that one dice of a pair of dice shows 4 if the sum of the two is greater than or equal to 9

OpenStudy (anonymous):

i calculated the probability to be 22/36 or 11/18

OpenStudy (dumbcow):

i get 1/9 * 36/10 = 2/5

OpenStudy (anonymous):

how did u get 1/9 and 36/10

OpenStudy (anonymous):

i know the only two possible choices were 4 and 5 to make the sum equal 9 or 10

OpenStudy (anonymous):

i meant 5 and 6

OpenStudy (dumbcow):

conditional probability: P(you have 4 AND sum >=9) = 4/36 P(sum >=9) = 10/36

OpenStudy (anonymous):

10/36?

OpenStudy (dumbcow):

4 ways of making 9 3 ways of making 10 2 ways of making 11 1 way of making 12

OpenStudy (anonymous):

ok. i get it now! how did u get 36? 6.6 because a dice has 6 faces

OpenStudy (dumbcow):

right , 6*6 = 36 possible combinations of rolling 2 dice

OpenStudy (anonymous):

i don't understand the question itself. it says that one dice should show 4

OpenStudy (anonymous):

so if one dice has to show 4 the possibility=1/6 for that dice. and not 4 is 5/6

OpenStudy (anonymous):

then for the other dice is has to either be 5 or 6. because 4+5=9 and 4+6=10

OpenStudy (anonymous):

u can't add anything to 4 to make it 11 or 12 with the other dice because a dice has 1-6

OpenStudy (dumbcow):

given sum>= 9, there are 10 possible ways the dice can be arranged right? then of those 10 only four have a 4 showing. 4,5 5,4 4,6 6,4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok i got it now! ways that the combination of both number is greater than or equal to 9 is 10.

OpenStudy (anonymous):

not including the fact that one has to be 4

OpenStudy (anonymous):

hence, the combination of numbers than can make the sum >= 9 is 10/36 whereas when one dice has 4 and the sum is >= is only 4 ways

OpenStudy (anonymous):

thanks for clearing it up for me :)

OpenStudy (dumbcow):

no problem

OpenStudy (anonymous):

one last question

OpenStudy (anonymous):

@dumbcow you had 4/36*36/10? why did u invert it? did u divide both probabilities?

OpenStudy (dumbcow):

comes from formula when using conditional probability \[P(A|B) = \frac{P(A and B)}{P(B)}\]

OpenStudy (dumbcow):

A -> one dice is 4 B -> sum >= 9

OpenStudy (anonymous):

oh ok! i forgot about that formula -_- thank you!

OpenStudy (anonymous):

thank u :)